# Devil is in Details: Locality-Aware 3D Abdominal CT Volume Generation for Self-Supervised Organ Segmentation

### v1

@@ 1. Introduction
Figure 2. Overview of Locality-Aware Diffusion (Lad).
@@ 3.1. Overview
As a crucial thread, the predicted mask of abdominal organ like the pancreas from a well-trained universal segmentation model, UniverSeg (Butoi et al., 2023) $f_{u}$, localizes the regions of anatomical structures and serves as a priori where we can dig into abdominal details. We denote the mask of the $i_{\text{th}}$ abdominal CT volume $x^{i}$, as $y^{i}$, $i=0,1,...,|D_{tr}|$, $D_{tr}$ means the trainset of abdominal CT volumes. The precision of the mask is insignificant. In scenarios where the prediction aligns closely with ground truth, detailed features of the abdominal organ can be effectively extracted by the mask. Conversely, inaccuracies in the prediction, indicating misalignment with the ground truth, present challenges for the segmentation model in identifying the correct regions. Therefore, enhancing precision in these areas is crucial for providing discriminative information.
@@ 3.2. Locality Refinement
Given the computational demands, the diffusion model operates within the latent space of vector quantized autoencoders (VQ-GAN) (Esser et al., 2021). The quality of the latent space directly impacts the fidelity of images generated by the diffusion model. To ensure the latent space fully captures the features of the original data, we define a loss function $\mathcal{L}_{\textit{glo}}$ for VQ-GAN, following (Khader et al., 2023):
@@
Apparently, all of these loss functions prioritize global reconstruction quality, ignoring the importance of detail reconstruction quality, particularly in anatomical structures. The significance of detail reconstruction varies across regions, where accurate depiction of anatomical structures holds paramount importance. Even if background reconstruction is excellent, blurred details of anatomical structures significantly hinder subsequent generation tasks. Thanks to the predicted mask $y^{i}$, we can localize the specific areas within volume $x^{i}$ that require attention. By cropping volume $x^{i}$ into a sub-volume $x_{sub}^{i}$ and performing the same operation on the reconstructed volume $\widehat{x^{i}}$, we obtain $\widehat{x_{sub}^{i}}$. We now introduce the Locality Loss $\mathcal{L}_{loc}$. This loss function is tailored to direct the reconstruction model’s focus towards the sub-volume $x_{sub}^{i}$ and the corresponding reconstructed sub-volume $\widehat{x_{sub}^{i}}$, leveraging the predicted mask $y^{i}$ to enhance the detail reconstruction of abdominal structures and thereby improve the quality of the latent space. The Locality Loss is formulated as the $L_{1}$ distance between the sub-volume $x_{sub}^{i}$ and the reconstructed sub-volume $\widehat{x_{sub}^{i}}$:
(2) $\mathcal{L}_{\textit{loc}}=\|(x_{sub}^{i},\widehat{x_{sub}^{i}})\|_{1},$
@@ 3.3.2. Locality Condition Extraction
To entail condition $c^{i}$ with abdominal features, we first use the encoder of the former trained VQ-GAN $E_{vq}$ to encode mask $y^{i}$ into the latent space of abdominal CT volume $x^{i}$ and then pass it through two convolution layers in order to save computation memory. We call this extraction process Content Extractor $E_{con}$ and denote the output as content vector $c^{i}_{con}$. $c^{i}_{con}$ contains image-level information such as the size, shape and location of the abdominal organ. Since $E_{vq}$ is trained on abdominal CT volume trainset $D_{tr}$, content vector $c^{i}_{con}$ represents encoding mask $y^{i}$ into an image-level latent space, without making the most of the structure information of mask label. Besides, the content vector $c^{i}_{con}$ represents the global features of the entire mask. Apparently, the significance of the background cannot match up to the foreground’s.
@@ 4.3.1. Synthesis for Training
We substitute synthetic data for real data in the segmentation model training to evaluate the genuine impact of synthetic data on downstream feature learning tasks, without incorporating any real data in the process. As depicted in Figure LABEL:fig:seg(a),fig:seg_ts(a), the segmentation model trained on synthetic volumes from Lad outperforms the one trained on synthetic volumes from Medical Diffusion in mean dice scores of four abdominal organs. As evidenced by higher dice scores for small organs such as the pancreas and spleen, our method’s emphasis on granularity significantly enhances the delineation of anatomical structures’ details. Despite prioritizing local features, our method achieves comparable performance to Medical Diffusion on larger organs such as the liver and kidney.
@@ 4.3.2. Synthesis for Augmentation
We incorporate synthetic volumes, which are approximately 20% the size of the training set, as a form of data augmentation for the training. Figure LABEL:fig:seg(b),fig:seg_ts(b) illustrates that our method yields the highest mean dice scores, thereby facilitating effective learning of visual representations. Notably, our method demonstrates particular strength in addressing the challenges with small organs, effectively bridging the performance gap observed when compared to Medical Diffusion.

### v2

@@ 1. Introduction
Figure 2. Overview of Locality-Aware Diffusion (Lad). Lad consists of three phases: (a) Latent space construction. We introduce a locality refinement module into the VQ-GAN training. Refinement module uses the Locality Loss $\mathcal{L}_{loc}$ to facilitate VQ-GAN to learn more details of anatomical structures in the low-dimensional space. (b) Diffusion fitting in latent space. We introduce a locality condition extraction module into the diffusion training. The diffusion model incorporates a locality condition extracted by Condition Extractor $E_{c}$ to generate volumes. (c) Sampling in latent space. We introduce a locality condition augmentation module into the diffusion sampling. In the augmentation module, diverse conditions are extracted from the augmented maskset and used to generate massive volumes. All three phases use the masks output by a priori extraction module. Masks are predicted by a well-trained universal segmentation model UniverSeg, with masks considered as priori.
@@ 3.1. Overview
As a crucial thread, the predicted mask of abdominal organ like the pancreas from a well-trained universal segmentation model, UniverSeg (Butoi et al., 2023) $f_{u}$, localizes the regions of anatomical structures and serves as a priori where we can dig into abdominal details. We denote the mask of the $i_{\text{th}}$ abdominal CT volume $x^{i}$, as $y^{i}$, $i=1,2,...,|\mathcal{D}_{tr}|$, $\mathcal{D}_{tr}$ means the trainset of abdominal CT volumes. The precision of the mask is insignificant. In scenarios where the prediction aligns closely with ground truth, detailed features of the abdominal organ can be effectively extracted by the mask. Conversely, inaccuracies in the prediction, indicating misalignment with the ground truth, present challenges for the segmentation model in identifying the correct regions. Therefore, enhancing precision in these areas is crucial for providing discriminative information.
@@ 3.2. Locality Refinement
Given the computational demands, the diffusion model operates within the latent space of VQ-GAN (Esser et al., 2021). The quality of the latent space directly impacts the fidelity of images generated by the diffusion model. To ensure the latent space fully captures the features of the original data, we define a loss function $\mathcal{L}_{\textit{glo}}$ for VQ-GAN, following (Khader et al., 2023):
@@
Apparently, all of these loss functions prioritize global reconstruction quality, ignoring the importance of detail reconstruction quality, particularly in anatomical structures. The significance of detail reconstruction varies across regions, where accurate depiction of anatomical structures holds paramount importance. Even if background reconstruction is excellent, blurred details of anatomical structures significantly hinder subsequent generation tasks. Thanks to the predicted mask $y^{i}$, we can localize the specific areas within volume $x^{i}$ that require attention. By cropping volume $x^{i}$ into a sub-volume $x_{\text{sub}}^{i}$ and performing the same operation on the reconstructed volume $\hat{x}^{i}$, we obtain $\hat{x}_{\text{sub}}^{i}$. We now introduce the Locality Loss $\mathcal{L}_{loc}$. This loss function is tailored to direct the reconstruction model’s focus towards the sub-volume $x_{\text{sub}}^{i}$ and the corresponding reconstructed sub-volume $\hat{x}_{\text{sub}}^{i}$, leveraging the predicted mask $y^{i}$ to enhance the detail reconstruction of abdominal structures and thereby improve the quality of the latent space. The Locality Loss is formulated as the $L_{1}$ distance between the sub-volume $x_{\text{sub}}^{i}$ and the reconstructed sub-volume $\hat{x}_{\text{sub}}^{i}$:
(2) $\mathcal{L}_{\textit{loc}}=\frac{\sum_{i=1}^{|\mathcal{D}_{tr}|}|x_{\text{sub}}^{i}-\hat{x}_{\text{sub}}^{i}|}{|\mathcal{D}_{tr}|}$
@@ 3.3.2. Locality Condition Extraction
To entail condition $c^{i}$ with abdominal features, we first use the encoder of the former trained VQ-GAN $E_{vq}$ to encode mask $y^{i}$ into the latent space of abdominal CT volume $x^{i}$ and then pass it through two convolution layers in order to save computation memory. We call this extraction process Content Extractor $E_{con}$ and denote the output as content vector $c^{i}_{con}$. $c^{i}_{con}$ contains image-level information such as the size, shape and location of the abdominal organ. Since $E_{vq}$ is trained on abdominal CT volume trainset $\mathcal{D}_{tr}$, content vector $c^{i}_{con}$ represents encoding mask $y^{i}$ into an image-level latent space, without making the most of the structure information of mask label. Besides, the content vector $c^{i}_{con}$ represents the global features of the entire mask. Apparently, the significance of the background cannot match up to the foreground’s.
@@ 4.3.1. Synthesis for Training
We substitute synthetic data for real data in the segmentation model training to evaluate the genuine impact of synthetic data on downstream feature learning tasks, without incorporating any real data in the process. As depicted in Figure 6(a) and 6(b), the segmentation model trained on synthetic volumes from Lad outperforms the one trained on synthetic volumes from Medical Diffusion in mean dice scores of four abdominal organs. As evidenced by higher dice scores for small organs such as the pancreas and spleen, our method’s emphasis on granularity significantly enhances the delineation of anatomical structures’ details. Despite prioritizing local features, our method achieves comparable performance to Medical Diffusion on larger organs such as the liver and kidney.
@@ 4.3.2. Synthesis for Augmentation
We incorporate synthetic volumes, which are approximately 20% the size of the training set, as a form of data augmentation for the training. Figure 6(c) and 6(d) illustrate that our method yields the highest mean dice scores, thereby facilitating effective learning of visual representations. Notably, our method demonstrates particular strength in addressing the challenges with small organs, effectively bridging the performance gap observed when compared to Medical Diffusion.